OpenStudy (anonymous):

The sides of a triangle are 80cm , 100 cm and 140 cm. Determine the radius of the inscribed circle

OpenStudy (anonymous):

Is it 40cm? Do u've the answer that u can check?

OpenStudy (anonymous):

r = 66.144 cm http://answers.yahoo.com/question/index?qid=20090217015757AAaWiPl

OpenStudy (anonymous):

The answer says that A1 = 1/2 ( 80 ) r = 40r A2 = 1/2 ( 100 ) r = 50r A3 = 1/2 (120) r = 60r but shouldn't A3 be 1/2(140)r? I'm not sure ab that~~

OpenStudy (anonymous):

The answer is not 40cm, sorry I hypothetically assumed two of the lengths are parallel~~I'm working on it~~

OpenStudy (mathmale):

I'd suggest making a good drawing of this triangle and the inscribed circle, and then drawing a radius from the center of the circle to each of the 3 sides. Note that each radius must be perpendicular to the corresponding side. Now draw a line from each vertex of the given triangle to the center of the circle. We'll end up with 6 smaller right triangles inside the given triangle. This info should be sufficient to enable finding the 3 interior angles of the given triangle, using the law of cosines. Anyone care to create and share such a drawing of this situation?

OpenStudy (anonymous):

Actually you can now solve the question as long as you have an example, however, I tried so I want to tell you what I did anyway. I convert the 80, 100 and 140 to 4,5,7 just to make things easier. because of the circle in side the triangle, you can have three triangles that have the same hight but different bases (4,5 and 7) A1=4r/2 A2= 7r/s A3=5r/2 then the total area of the triangle is 8r. Then you have the formula that p=(a+b+c)/2 which in this case equal to 8 and then area=squroot(8-4)(8-7)(8-5)8 which equals to 4times squrot 6. Then u equal the 8r to the 4squrot6 and solve it. I'm not quite sure but the numbers but the method I think is right. remember to convert it back after the calculation, u need to time them by 20. Good luck with that.

OpenStudy (anonymous):

actually the answer is 24.49 cm

OpenStudy (mathmale):

@ eranronald: How about dropping a few clues regarding how you arrived at your assumed solution? What methods did you use? Possibly the Law of Sines? the Law of Cosines? the concept of "median"? ??